# FOCUS on Contamination: A Geospatial Deep Learning Framework with a Noise-Aware Loss for Surface Water PFAS Prediction

### v3

@@
Abstract
Per- and polyfluoroalkyl substances (PFAS), chemicals found in products like non-stick cookware, are unfortunately persistent environmental pollutants with severe health risks. Accurately mapping PFAS contamination is crucial for guiding targeted remediation efforts and protecting public and environmental health, yet detection across large regions remains challenging due to the cost of testing and the difficulty of simulating their spread. In this work, we introduce FOCUS, a geospatial deep learning framework with a label noise-aware loss function, to predict PFAS contamination in surface water over large regions. By integrating hydrological flow data, land cover information, and proximity to known PFAS sources, our approach leverages both spatial and environmental context to improve prediction accuracy. We evaluate the performance of our approach through extensive ablation studies, robustness analysis, real-world validation, and comparative analyses against baselines like sparse segmentation, as well as existing scientific methods, including Kriging and pollutant transport simulations. Results and expert feedback highlight our framework’s potential for scalable PFAS monitoring.
Introduction
@@
PFAS or per- and polyfluoroalkyl substances are persistent “forever chemicals” widely used in industrial and consumer products such as non-stick cookware, waterproof textiles, and firefighting foams (National Institute of Environmental Health Sciences n.d.). Unfortunately, these substances resist degradation and accumulate in water (Langenbach and Wilson 2021; Schroeder, Bond, and Foley 2021), soil (Crone et al. 2019), and living organisms (Environmental Working Group 2023b), with 97% of Americans exhibiting detectable levels in their blood (CDC 2024). Such pervasive exposure is linked to severe health risks including cancers, liver damage, and developmental disorders (Manz 2024). With the high cost of measuring PFAS concentrations in the environment, which leads to sparse ground truth data (Fig. 1), we are left uncertain about targeted remediation to best protect health.
Figure 1: Illustration of limited PFAS contamination data (U.S. EPA, OW 2015)
@@
Figure 2: Left: Classical ML techniques aggregate surrounding pixel data to predict contamination at specific points (depicted as x, y coordinates); Right: DL methods process raster images directly to generate dense PFAS contamination maps in a single pass.
@@
Artificial intelligence (AI) presents a promising avenue for addressing complex health and environmental issues, including identifying PFAS hotspots. In various domains, AI has proven effective at automating labor-intensive tasks and synthesizing domain expertise, whether in land cover mapping (Robinson et al. 2020), agricultural optimization (Kerner et al. 2024), or disease prediction (Bondi-Kelly et al. 2023). In the domain of PFAS contamination prediction, scientists have predominantly applied machine learning (ML) models such as random forests (Breiman 2001) and XGBoost (Chen and Guestrin 2016). For example, (DeLuca et al. 2023) uses a random forest model to predict PFAS contamination via a tabular approach that aggregates environmental features within a 5 km buffer around sample points (Fig. 2). Building on these efforts, we frame PFAS prediction as a geospatial deep learning (DL) task. Our approach can take satellite image formats (i.e., rasters) as input directly, preserving spatial dependencies, reducing the need for extensive feature engineering and aggregation, and leading to more efficiency for ongoing monitoring efforts.
@@
  • We propose FOCUS (Framework for EnvirOnmental Contamination with Uncertainty Scaling), a framework with a geospatial DL model for segmentation of PFAS contamination using image formats (rasters) directly.
  • We account for the limited labeled data by developing a novel noise-aware loss, leveraging domain expertise on PFAS spread to fill data gaps with confidence weighting.
  • We benchmark our sparse segmentation framework against other AI models (Hua et al. 2022), tabular ML approaches (DeLuca et al. 2023), and popular methods like pollutant transport simulations (SWAT n.d.) and Kriging (GISGeography 2017), demonstrating our framework’s scalability and accuracy for large-scale PFAS prediction.
  • Through ablation studies, we assess our FOCUS loss function and hyperparameter selection.
We develop this framework together with an environmental nonprofit organization, academic researchers specializing in water quality models and environmental chemistry, and an advisory group with representatives from other nonprofits and agencies, ensuring that our approach is grounded in the latest expertise and real-world environmental concerns.
Related Work
@@
The measurement of PFAS relies on laboratory-based analysis of water, fish tissue, and human blood samples with techniques such as liquid chromatography-tandem mass spectrometry (LC-MS) (Shoemaker and Tettenhorst 2020). PFAS monitoring efforts have made use of such measurement techniques and beyond. For example, states such as Colorado have implemented water testing, treatment grants, and regulatory measures since 2016 (Colorado Department of Public Health and Environment n.d.). In addition, there are community-based water sampling initiatives such as volunteer sampling by the Sierra Club (Sierra Club n.d.) and focused efforts by organizations like the Ecology Center (EcoCenter n.d.). A recent study further underscores the importance of these efforts, revealing associations between PFAS contamination in drinking water and higher COVID-19 mortality rates in the U.S. (Liddie et al. 2024). Resources like the PFAS-Tox Database (Pelch et al. 2022) also assist communities in monitoring PFAS. While highly accurate and important, laboratory-based PFAS measurements are costly, time-intensive, and difficult to scale spatially (Doudrick 2024).
@@
Various modeling approaches have been explored to predict contaminant distribution. Hydrological models such as SWAT (SWAT n.d.) and MODFLOW (Torcoletti 2012) simulate pollutant transport using environmental parameters like flow rates, land use, and weather conditions. While effective, these models require extensive data and are computationally expensive at large scales (Zhi et al. 2024). Kriging, a geostatistical interpolation technique, has been widely used for mapping pollutants, including in soil (Largueche 2006) and groundwater (Singh and Verma 2019). Though useful in data-sparse environments, Kriging relies on simplifying assumptions such as spatial continuity and stationarity (GISGeography 2017) that may not hold for complex contaminants like PFAS, whose transport dynamics and uncertainties are more intricate (see section Results). Complementing these efforts, Salvatore et al. (Salvatore et al. 2022) proposed a “presumptive contamination” model that estimates PFAS presence based solely on proximity to known PFAS-using or -producing facilities. We aim to contribute to monitoring and modeling efforts by providing scalable, data-driven PFAS predictions.
AI Approaches for PFAS Prediction.
@@
Recent advances in AI and geospatial modeling have opened new avenues for tackling the scalability challenges of PFAS detection. For example, DeLuca et al. (DeLuca et al. 2023) used random forests to predict PFAS contamination in fish tissue in the Columbia River Basin, leveraging geospatial data from industrial and military facilities. On a national scale, the USGS applied XGBoost to forecast PFAS occurrence in groundwater (Tokranov et al. 2023) using PFAS sources, groundwater recharge, etc. Such approaches require extensive manual feature engineering, such as aggregating data spatially before prediction is possible (Fig. 2). In addition to being resource-intensive (see section Results), this preprocessing loses spatial context.
Geospatial Models.
@@
Recent advances in geospatial deep learning have introduced foundation models trained on large-scale satellite datasets for improved spatial prediction. Models like Prithvi (Blumenfeld 2023) and SatMAE (Cong et al. 2023) use masked autoencoders (MAE) to learn generalizable spatial representations in a semi-supervised manner, leveraging both labeled and unlabeled data. Unlike traditional ML approaches that require extensive feature engineering (e.g., Fig. 2), these models process raster (image) data directly, preserving spatial dependencies. However, many of these approaches assume that key information is directly observable in satellite imagery, which is not the case for PFAS pollution (see section Results). We therefore adapt Prithvi to use preprocessed, proxy features, such as land cover.
Addressing Data Scarcity.
@@
In environmental data analysis, limited ground truth often necessitates pseudo- or augmented labels, which can introduce noise and lead to overfitting (Safonova et al. 2023). Various strategies have been proposed to mitigate these issues, including label smoothing and robust loss functions (e.g., Huber loss (Gokcesu and Gokcesu 2021), Generalized Cross-Entropy (Zhang and Sabuncu 2018), bootstrapping loss (Reed et al. 2015)), and focal loss (Lin et al. 2018)). Among these, focal loss is particularly effective for addressing class imbalance by down-weighting well-classified examples and focusing the model’s learning on hard, uncertain samples. Weakly supervised methods like the FESTA loss (Hua et al. 2022) further aim to capture spatial and feature relationships. FESTA is specifically designed for segmentation of remote sensing images using sparse annotations assuming spatial continuity, which may not hold for PFAS (see section Results). Furthermore, our incomplete understanding of PFAS-specific hydrological processes hinders the application of both transfer and active learning approaches, making it difficult to estimate uncertainty and find reasonable proxies (Guelfo et al. 2021; Russo et al. 2020). These challenges underscore the need for methods that integrate domain-specific knowledge directly into the training process.
Methodology
This study aims to predict PFAS presence to inform environmental decision-making. In particular, we perform segmentation on surface water area, classifying them as having PFAS concentrations above or below established health advisory thresholds using geospatial data. We next describe FOCUS, including the data preparation pipeline, and the modeling and evaluation approaches.
Figure 3: Overview of dataset curation pipeline
Dataset
@@
Ground Truth Samples
We use a combination of datasets and health advisories from government agencies in the United States, for both water and fish tissue samples. These are summarized in Table 1 (details in supplement).
@@
Fish Train/Test EPA (U.S. EPA, OW 2015; US EPA 2024) 866
Fish Validation MPART (Michigan Department of Environment, Great Lakes, and Energy 2025) 114
Water Train/Test MPART (Michigan Department of Environment, Great Lakes, and Energy 2023) 293
@@
Table 1: Datasets used for training, testing, & real-world validation.
We utilize both fish and water data, as they provide snapshots of PFAS contamination with different contexts. For example, fish accumulate PFAS over time, serving as longer-term indicators of water contamination and providing insights into potential exposure risks to anglers (Environmental Working Group 2023a), whereas water sampling provides a measure of PFAS contamination at a specific place and time. Our analysis primarily uses fish data due to the nationally representative extent.
@@
Furthermore, PFAS concentrations are measured as continuous values in these samples, with individual measurements for different species of PFAS. We binarize PFAS concentration measurements into above (1) or below (0) health advisory thresholds, such as the EPA’s Fish and Shellfish Advisory Program health thresholds (U.S. EPA 2024). In particular, we use the cumulative metric called the Hazard Index, which assesses combined risk from individual PFAS compounds by summing their concentration-to-threshold ratios (U.S. Environmental Protection Agency 2023). Note that for fish train/test data, this yields 775 above-threshold (89.5%) and 91 below-threshold (10.5%) samples, reflecting a critical real-world imbalance. For completeness, we also experimented with an alternative labeling scheme using three classes (see supplement).
Raster Image Generation
@@
To spatially integrate the data for training and testing, multi-channel raster images with were generated around each sample point. Each image is a $P\times P$ pixels patch with 30-meter resolution, centered on the geographic coordinates of a sample point, where the optimal value of $P$ is determined in Section 3.3. Once trained, the model can make predictions at any location simply by extracting and processing the corresponding raster, which is done for real-world validation.
@@
These raster images integrate several key features. First, we incorporate readily available rasters such as the National Land Cover Database (NLCD) (U.S. Geological Survey 2024), and flow direction rasters that capture hydrological connectivity using the D8 algorithm in ArcGIS Pro (Esri 2024). In addition, we include discharger location data obtained from the U.S. EPA Enforcement and Compliance History Online (ECHO) (U.S. Environmental Protection Agency n.d.) and convert them into distance rasters using a distance transform (SciPy Developers 2019) (Fig. 4).
@@
Figure 4: Example raster channels: (left) land cover, (center) distances from chemical manufacturing industries, and (right) flow direction.
Ground Truth Masks
@@
Our goal is to perform segmentation to predict PFAS contamination in surface water, but we have sparse point-level contamination measurements rather than dense labels across rasters. To generate dense training labels, every surface water pixel in a patch is labeled according to the point-level sample measurement in that patch (i.e., if the sample point indicates PFAS above the safety thresholds, all surface water pixels are labeled 1; if below, they are labeled 0). Non-surface water pixels are assigned a value of 2. This approach extends point-level data across the patch, assuming that nearby areas have similar contamination.
Noise Masks
This assumption may generate inaccurate labels. To quantify the confidence that a pixel’s label (either 0 or 1) is correct after expansion, we generate a domain expert-informed probability of the label correctness for each surface water pixel $i$:
$\,+\alpha_{4}\cdot p_{\text{downstream}}$ (1)
  • Proximity to PFAS Dischargers ($p_{\text{dischargers}}$) Proximity to known PFAS dischargers is a strong indicator of contamination risk, as shown in scientific analyses (Salvatore et al. 2022). If a pixel’s label is 1, we apply an exponential decay function of distance, assigning higher probabilities to pixels closer to known dischargers. For pixels labeled 0, we invert this logic so that being near a discharger lowers confidence in the 0 label.
  • Land Cover Type ($p_{\text{landcover}}$) For pixels labeled 1, surrounding urban or built-up areas received higher probabilities, consistent with research linking PFAS to industrial zones (Dimitrakopoulou et al. 2024). Conversely, for 0-labeled pixels, surrounding undeveloped or forested land cover increased confidence in low contamination.
  • Proximity to Other Sample Points ($p_{\text{sample\_dist}}$) Pixels closer to ground truth sample points were assigned higher probabilities using an exponential decay function based on distance to each sample point in the patch.
  • Downstream Flow ($p_{\text{downstream}}$) PFAS may also travel downstream in aquatic systems, making hydrologic flow another critical factor in contamination spread (Pétré et al. 2022). If a pixel lies downstream of a sample point labeled 1, its probability for label 1 increases, and similarly for label 0. This flow direction channel was generated using ArcGIS Pro (ESRI n.d.) based on digital elevation models (DEMs) (Earth Science Data Systems 2024).

### v4

# FOCUS on Contamination: Hydrology-Informed Noise-Aware Learning for Geospatial PFAS Mapping

@@
Abstract.
Per- and polyfluoroalkyl substances (PFAS) are persistent environmental contaminants with significant public health impacts, yet large-scale monitoring remains severely limited due to the high cost and logistical challenges of field sampling. The lack of samples leads to difficulty simulating their spread with physical models and limited scientific understanding of PFAS transport in surface waters. Yet, rich geospatial and satellite-derived data describing land cover, hydrology, and industrial activity are widely available. We introduce FOCUS, a geospatial deep learning framework for PFAS contamination mapping that integrates sparse PFAS observations with large-scale environmental context, including priors derived from hydrological connectivity, land cover, source proximity, and sampling distance. These priors are integrated into a principled, noise-aware loss, yielding a robust training objective under sparse labels. Across extensive ablations, robustness analyses, and real-world validation, FOCUS consistently outperforms baselines including sparse segmentation, Kriging, and pollutant transport simulations, while preserving spatial coherence and scalability over large regions. Our results demonstrate how AI can support environmental science by providing screening-level risk maps that prioritize follow-up sampling and help connect potential sources to surface-water contamination patterns in the absence of complete physical models.
journalyear: 2026doi: XXXXXXX.XXXXXXXcopyright: noneconference: Make sure to enter the correct conference title from your rights confirmation email; Washington, DC; copyright: noneconference: ; ; $c$s: Computing methodologies Machine learningccs: Computing methodologies Learning with noisy labelsccs: Applied computing Environmental sciences
1. Introduction
@@
PFAS, or per- and polyfluoroalkyl substances, are persistent “forever chemicals” widely used in industrial and consumer products such as non-stick cookware, waterproof textiles, and firefighting foams (National Institute of Environmental Health Sciences, nd). Due to their extreme chemical stability, PFAS resist degradation and are now pervasive across environmental compartments, accumulating in surface and groundwater (Langenbach and Wilson, 2021; Schroeder et al., 2021), soils (Crone et al., 2019), and living organisms (Environmental Working Group, 2023b). This widespread environmental presence has direct public health consequences, with 97% of Americans exhibiting detectable levels in their blood (CDC, 2024) and exposure linked to cancers, liver damage, and developmental disorders (Manz, 2024). However, measuring PFAS concentrations remains expensive—often costing hundreds of U.S. dollars per sample—resulting in sparse and unevenly distributed ground-truth data (Fig. 1). This data scarcity leaves substantial uncertainty in identifying contamination hotspots and prioritizing targeted remediation to best protect public health.
Figure 1. Illustration of limited PFAS contamination data (U.S. EPA, OW, 2015)
@@
Figure 2. Left: Classical ML techniques aggregate surrounding pixel data to predict contamination at specific points (depicted as x, y coordinates); Right: DL methods process raster images directly to generate dense PFAS contamination maps in a single pass.
@@
Artificial intelligence (AI) presents a promising avenue for addressing complex health and environmental issues, including identifying PFAS hotspots. In various domains, AI has proven effective at automating labor-intensive tasks and synthesizing domain expertise, whether in land cover mapping (Robinson et al., 2020), agricultural optimization (Kerner et al., 2024), disease prediction (Bondi-Kelly et al., 2023), or environmental and Earth-system inference tasks (Reichstein et al., 2019; Christin et al., 2019). In the domain of PFAS contamination prediction, however, prior work has largely relied on tabular ML models such as random forests (Breiman, 2001) and XGBoost (Chen and Guestrin, 2016), which aggregate environmental features around sparse sample points and discard spatial structure. For example, (DeLuca et al., 2023) uses a random forest model with features aggregated within a 5 km buffer around each sample (Fig. 2). Building on these efforts, we instead frame PFAS prediction as a geospatial deep learning (DL) task that operates directly on raster inputs, preserving spatial dependencies while reducing the need for manual feature engineering and aggregation.
@@
  • We propose FOCUS, a principled noise-aware geospatial learning framework comprising a hydrology informed noise-aware objective that combines hydrological priors into a focal-style loss to train robust segmentation models from sparse, noisy PFAS measurements.
  • We formalize PFAS surface-water mapping as learning under structured, spatially dependent label noise induced by expanding point measurements to pixel-level supervision.
  • We derive a theoretical result showing that FOCUS optimizes a valid surrogate objective under asymmetric, pixel-wise label noise, providing a principled justification for the proposed training loss.
  • We construct pixel-wise confidence weights from hydrological flow, industrial discharge proximity, land cover, and sampling distance, linking environmental processes to robust learning.
  • We evaluate on real PFAS field data across the United States and benchmark against geostatistical methods, pollutant transport simulations, and tabular ML baselines, demonstrating improved predictive performance and scalability.
Domain co-design and motivation. PFAS mapping poses domain-specific challenges that limit both traditional modeling and off-the-shelf ML: sparse and biased measurements, expensive and uncertain physical simulations, and heavy reliance on manual feature engineering and spatial aggregation. We developed FOCUS in close collaboration with environmental scientists from the Environmental Working Group, academic researchers specializing in water quality models and environmental chemistry, and an advisory group with representatives from other nonprofits and agencies such as the Huron River Watershed Council (HRWC), who helped shape both the model inputs (e.g., hydrology, land cover, source proximity) and the learning objective to reflect real scientific uncertainty and regulatory priorities. This co-design process informed our choice to operate directly on geospatial rasters, and to encode physically motivated correctness priors.
Reproducibility. Code, processed datasets, and scripts to reproduce the FOCUS experiments are available in the accompanying repository  [^1] Code available at https://anonymous.4open.science/r/FOCUS_KDD-D184/README.md..
2. Related Work
@@
Current Ways to Measure and Predict PFAS. The measurement of PFAS relies on laboratory-based analysis of water, fish tissue, and human blood samples with techniques such as liquid chromatography-tandem mass spectrometry (LC-MS) (Shoemaker and Tettenhorst, 2020). PFAS monitoring efforts have made use of such measurement techniques and beyond. For example, states such as Colorado have implemented water testing, treatment grants, and regulatory measures since 2016 (Colorado Department of Public Health and Environment, nd). In addition, there are community-based water sampling initiatives such as volunteer sampling by the Sierra Club (Sierra Club, nd) and focused efforts by organizations like the Ecology Center (EcoCenter, nd). A recent study further underscores the importance of these efforts, revealing associations between PFAS contamination in drinking water and higher COVID-19 mortality rates in the U.S. (Liddie et al., 2024). Resources like the PFAS-Tox Database (Pelch et al., 2022) also assist communities in monitoring PFAS. While highly accurate and important, laboratory-based PFAS measurements are costly, time-intensive, and difficult to scale spatially (Doudrick, 2024).
@@
Various modeling approaches have been explored to predict contaminant distribution. Hydrological models such as SWAT (SWAT, nd) and MODFLOW (Torcoletti, 2012) simulate pollutant transport using environmental parameters like flow rates, land use, and weather conditions. While effective, these models require extensive data and are computationally expensive at large scales (Zhi et al., 2024). Kriging, a geostatistical interpolation technique, has been widely used for mapping pollutants, including in soil (Largueche, 2006) and groundwater (Singh and Verma, 2019). Though useful in data-sparse environments, Kriging relies on simplifying assumptions such as spatial continuity and stationarity (GISGeography, 2017) that may not hold for complex contaminants like PFAS, whose transport dynamics and uncertainties are more intricate (see Section 5). Complementing these efforts, Salvatore et al. (Salvatore et al., 2022) proposed a “presumptive contamination” model that estimates PFAS presence based solely on proximity to known PFAS-using or -producing facilities. We aim to contribute to monitoring and modeling efforts by providing scalable, data-driven PFAS predictions.
@@
AI Approaches for PFAS Prediction. Recent advances in AI and geospatial modeling have opened new avenues for tackling the scalability challenges of PFAS detection. For example, DeLuca et al. (DeLuca et al., 2023) used random forests to predict PFAS contamination in fish tissue in the Columbia River Basin, leveraging geospatial data from industrial and military facilities. On a national scale, the USGS applied XGBoost to forecast PFAS occurrence in groundwater (Tokranov et al., 2023) using PFAS sources, groundwater recharge, etc. Such approaches require extensive manual feature engineering, such as aggregating data spatially before prediction is possible (Fig. 2). In addition to being resource-intensive (see Section 5), this preprocessing loses spatial context.
@@
Geospatial Models. Recent advances in geospatial deep learning have introduced foundation models trained on large-scale satellite datasets for improved spatial prediction. Neural operators and physics-informed models target PDE systems (Li et al., 2020; Karniadakis et al., 2021); however, PFAS transport lacks reliable governing physics and dense simulations, motivating a noise-aware, data-driven geospatial approach. Models like Prithvi (Blumenfeld, 2023) and SatMAE (Cong et al., 2023) use masked autoencoders (MAE) to learn generalizable spatial representations in a semi-supervised manner, leveraging both labeled and unlabeled data. Unlike traditional ML approaches that require extensive feature engineering (e.g., Fig. 2), these models process raster (image) data directly, preserving spatial dependencies. However, many of these approaches assume that key information is directly observable in satellite imagery, which is not the case for PFAS pollution (see Section 5). We therefore adapt Prithvi to use preprocessed, proxy features, such as land cover.
@@
Addressing Data Scarcity. In environmental data analysis, limited ground truth often necessitates pseudo- or augmented labels, which can introduce noise and lead to overfitting (Safonova et al., 2023). Various strategies have been proposed to mitigate these issues, including label smoothing and robust loss functions (e.g., Huber loss (Gokcesu and Gokcesu, 2021), Generalized Cross-Entropy (Zhang and Sabuncu, 2018), bootstrapping loss (Reed et al., 2015)), and focal loss (Lin et al., 2018)). Among these, focal loss is particularly effective for addressing class imbalance by down-weighting well-classified examples and focusing the model’s learning on hard, uncertain samples. Weakly supervised methods like the FESTA loss (Hua et al., 2022) further aim to capture spatial and feature relationships. FESTA is specifically designed for segmentation of remote sensing images using sparse annotations assuming spatial continuity, which may not hold for PFAS (see Section 5). Moreover, limited understanding of PFAS hydrological processes complicates the use of transfer and active learning, making it difficult to estimate uncertainty and identify reliable proxies (Guelfo et al., 2021; Russo et al., 2020). These challenges underscore the need for methods that integrate domain-specific knowledge directly into the training process.
Figure 3. Overview of dataset curation pipeline
3. Methodology
PFAS mapping presents unique challenges: labels exist only at sparse monitoring locations, yet contamination spreads via hydrological and land-use processes. To address this, we propose FOCUS as a geospatial learning framework with a noise-aware objective grounded in a latent contamination model and hydrology-informed pixel correctness priors. We perform segmentation on surface water area, classifying them as having PFAS concentrations above or below established health advisory thresholds using geospatial data.
3.1. Dataset
@@
3.1.1. Ground Truth Samples
We use a combination of datasets and health advisories from government agencies in the United States, for both water and fish tissue samples. These are summarized in Table 1 (details in appendix).
Table 1. Datasets used for training, testing, and real-world validation.
@@ 3.1.1. Ground Truth Samples
Fish Train/Test EPA (U.S. EPA, OW, 2015; US EPA, 2024) 866
Fish Validation MPART (Michigan Department of Environment, Great Lakes, and Energy, 2025) 114
Water Train/Test MPART (Michigan Department of Environment, Great Lakes, and Energy, 2023) 293
@@
We use both fish and water samples: fish reflect longer-term accumulation relevant to exposure risk, while water samples capture localized contamination at a given time (Environmental Working Group, 2023a). Our analysis primarily uses fish data due to the nationally representative extent.
@@
Furthermore, PFAS concentrations are measured as continuous values in these samples, with individual measurements for different species of PFAS. We binarize PFAS concentration measurements into above (1) or below (0) health advisory thresholds, such as the EPA’s Fish and Shellfish Advisory Program health thresholds (U.S. EPA, 2024). In particular, we use the cumulative metric called the Hazard Index, which assesses combined risk from individual PFAS compounds by summing their concentration-to-threshold ratios (U.S. Environmental Protection Agency, 2023). Note that for fish train/test data, this yields 775 above-threshold (89.5%) and 91 below-threshold (10.5%) samples, reflecting a critical real-world imbalance. For completeness, we also experimented with an alternative labeling scheme using three classes (see appendix).
Figure 4. Model overview: high-level representation of FOCUS
3.1.2. Raster Image Generation
@@
To spatially integrate the data for training and testing, multi-channel raster images were generated around each sample point (see Fig 5). Each image is a $P\times P$ pixels patch with 30-meter resolution, centered on the geographic coordinates of a sample point, where the optimal value of $P$ is determined in Section 4.3. Once trained, the model can make predictions at any location simply by extracting and processing the corresponding raster, which is done for real-world validation.
@@
These raster images integrate several key features. First, we incorporate readily available rasters such as the National Land Cover Database (NLCD) (U.S. Geological Survey, 2024), and flow direction rasters that capture hydrological connectivity using the D8 algorithm in ArcGIS Pro (Esri, 2024). In addition, we include discharger location data obtained from the U.S. EPA Enforcement and Compliance History Online (ECHO) (U.S. Environmental Protection Agency, nd) and convert them into distance rasters using a distance transform (SciPy Developers, 2019).
@@
Figure 5. Example raster channels: (left) land cover, (center) distances from chemical manufacturing industries, and (right) flow direction.
3.1.3. Ground Truth Masks
@@
Our goal is to perform segmentation to predict PFAS contamination in surface water, but we have sparse point-level contamination measurements rather than dense labels across rasters. To generate dense training labels, every surface water pixel in a patch is labeled according to the point-level sample measurement in that patch (i.e., if the sample point indicates PFAS above the safety thresholds, all surface water pixels are labeled 1; if below, they are labeled 0). Non-surface water pixels are assigned a value of 2. This approach extends point-level data across the patch, assuming that nearby areas may have similar contamination.
3.2. Problem Setup
Let $z_{i}\in\{0,1\}$ denote the latent contamination state at pixel $i$, and $y_{i}\in\{0,1\}$ be the observed label obtained by expanding point measurements across surface water (see Section 3.1.3). Given pixel features $x_{i}$ (land cover, hydrology, distance to facilities, etc.), the goal is to learn
(1) $\pi_{\theta}(x_{i})=\Pr\!\left(z_{i}=1\mid x_{i}\right)$
under sparse and noisy supervision.
Two challenges arise:
  1. (1)
    Spatially-structured label noise: contamination spreads directionally (downstream), is source-driven (industrial discharge), and varies by land use.
  2. (2)
    Regulatory sensitivity: false negatives pose public health risk, requiring calibrated uncertainty and high recall on positives.
3.3. Latent Noise Model
We assume per-pixel asymmetric flip noise $\eta_{i}<\frac{1}{2}$, where $\eta_{i}=\Pr(y_{i}\neq z_{i}\mid x_{i})$, motivated by laboratory detection limits and geographic label propagation. The noisy likelihood is:
(2) $\Pr\!\left(y_{i}=1\mid x_{i}\right)=\eta_{i}+(1-2\eta_{i})\,\pi_{\theta}(x_{i})$
Unlike classical noisy-label models that assume global or class-conditional noise rates (Natarajan et al., 2013; Goldberger and Ben-Reuven, 2017), we model $\eta_{i}$ as pixel-wise and spatially varying, reflecting physical measurement limits and label propagation rather than annotation randomness. While instance-dependent noise has been studied (e.g., Confident Learning (Northcutt et al., 2021)), here $\eta_{i}$ is driven by geophysical processes and spatial structure, not inferred solely from prediction statistics.
3.4. Physically-Informed Pixel Confidence $M_{i}$ Details
We estimate a correctness score $M_{i}\in[0,1]$ for each training pixel using environmental priors:
(3) $M_{i}=\alpha_{1}\,p_{\text{discharger}}+\alpha_{2}\,p_{\text{landcover}}+\alpha_{3}\,p_{\text{sample}}+\alpha_{4}\,p_{\text{downstream}}$
where:
  • $p_{\text{discharger}}$ decays with distance to PFAS facilities,
  • $p_{\text{landcover}}$ reflects industrial vs. natural land types,
  • $p_{\text{sample}}$ decays from monitoring points,
  • $p_{\text{downstream}}$ increases along hydrological flow.
These terms map environmental processes to pixelwise label correctness priors. Weights $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ are selected based on domain knowledge and experimental validation.